Find the terminal point P ( x , y ) on the unit circle determined by the given value of t:
If t=π/2, then x =
AND
y=

Answer by math_tutor2020(3816)


Answers: x = 0, y = 1

The radian angle pi/2 is equivalent to 90 degrees. This angle points directly north, which corresponds to the point (0,1)
I recommend you have a unit circle reference sheet with you. It could be a chart/table of values and/or an actual diagram of the unit circle. You can decide which is your preference.

Note how
cos(t) = cos(pi/2) = 0
sin(t) = sin(pi/2) = 1
Make sure your calculator is in radian mode

It turns out that,
x = cos(t)
y = sin(t)
So any (x,y) point on the unit circle can be described as (cos(t), sin(t))